Bvg of mass 5x1o ka thes into a car and eis squashed. C was m to This happened in o.ols what uos the Sorce os the bua on the car? wha+ was the torce of the car on the bua? Impulse? hange in velocity of the bu ation Torce
Answer #1

Impulse= change in momentum =average force * time interval

Impulse=F_{av}*Delta t=mDelta v

Delta v =v_{f}-v_{i}

Given m5 10-6kq

v_{i}=1m/s

v_{f}=0m/s

LV 0.01s

F_{av}*Delta t=mDelta v

Fav * 0.01s 5 * 10-6kg(0m/s-1m/s)

Far * 0.01 =-5 * 10-6

F5 104N (-ve sign shows force is opposite to direction of motion)

============

Force of bug on car will be equal to average force

ANSWER: F= 5 * 10-4M

Force of car on bug will be equal and opposite to force of bug on car

ANSWER: F= 5 * 10-4M

==================

Impulse=F_{av}*Delta t

Impulse=5*10^{-4}*0.01s

ANSWER: Impulse = 5 * 10-oNs

================================